Find the general solution of the differential equation

y²/2 + 2y exp(t) + (y+exp(t)) dy/dt = 0
1
Expert's answer
2019-08-25T12:45:19-0400

"\frac{y^2}{2}+2ye^t+(y+e^t)dy/dt=0"

"(\frac{y^2}{2}+2ye^t)dt+(y+e^t)dy=0"

"P(t,y)=\frac{y^2}{2}+2ye^t"

"Q(t,y)=y+e^t"

"\u2202P\/\u2202y=y+2e^t"

"\u2202Q\/\u2202t=e^t"

"∂P/∂y \not = ∂Q/∂t", so equation is not exact.

Let's find an integrating factor "\mu" :

"\xi=\frac{∂P/∂y - ∂Q/∂t}{Q}=\frac{y+2e^t-e^t}{y+e^t}=\frac{y+e^t}{y+e^t}=1"

"\xi" is a function of t only (not y), let it be denoted by "\xi (t)". Then

"\mu=e^{\int\xi(t)dt}=e^{\int dt}=e^t"

Differential equation with integrating factor:

"(\frac{y^2e^t}{2}+2ye^{2t})dt+(ye^t+e^{2t})dy=0"

"P_1(t,y)=\frac{y^2e^t}{2}+2ye^{2t}"

"Q_1(t,y)=ye^t+e^{2t}"

"\u2202P_1\/\u2202y=ye^t+2e^{2t}"

"\u2202Q_1\/\u2202t=ye^t+2e^{2t}"

"\u2202P_1\/\u2202y = \u2202Q_1\/\u2202t", so equation "(\frac{y^2e^t}{2}+2ye^{2t})dt+(ye^t+e^{2t})dy=0" is exact.

So, solution of this equation "f(t,y)" and "\u2202f(t,y)\/\u2202t=P_1(t,y)" and "\u2202f(t,y)\/\u2202y=Q_1(t,y)".

"f(t,y)=\int P_1(t,y)dt +\phi(y)"

"f(t,y)=\int (\frac{y^2e^t}{2}+2ye^{2t})dt +\phi(y)=\frac{y^2e^t}{2}+ye^{2t} +\phi(y)"

"∂f(t,y)/∂y=(\frac{y^2e^t}{2}+ye^{2t} +\phi(y))'=ye^{t}+e^{2t}+\phi'(y) =Q_1(t,y)"

"ye^{t}+e^{2t}+\phi'(y) =ye^t+e^{2t}"

"\phi'(y) =0"

"\phi(y) =C_1"

"f(t,y)=\frac{y^2e^t}{2}+ye^{2t} +C_1"

So, a solution of the differential equation is

"f(t,y)=C_2"

"\frac{y^2e^t}{2}+ye^{2t} +C_1=C_2,\, C=C_2-C_1,"

"\frac{y^2e^t}{2}+ye^{2t} =C"


Answer: "\frac{y^2e^t}{2}+ye^{2t} =C"